Consider K = 1 −1 1 −1 then we get K2 = 0 Does this hold for real numbers? Motivate.
1
Expert's answer
2021-10-06T17:42:35-0400

If ab=0, either a=0 or b=0

-Products of two non-zero numbers is always non-zero

But products of two non-zero matrices can be zero matrix


Using K given above


K="\begin{bmatrix}
 1&-1&;&1&-1 
 
\end{bmatrix}"


That is K"\begin{pmatrix}
 1 & -1 \\
 1 & -1
\end{pmatrix}"

K2=(K)(K)


="\begin{pmatrix}
 1 & -1 \\
 1& -1
\end{pmatrix}" "\begin{pmatrix}
 1 & -1\\
 1 & -1
\end{pmatrix}"


"\begin{pmatrix}
 1×1+-1×1 & 1×-1+-1×-1\\
 1×1+-1×1 & 1×-1+-1×-1
\end{pmatrix}"



="\begin{pmatrix}
 0 & 0\\
 0& 0
\end{pmatrix}"


=0 (Null matrix)